OpenStudy (anonymous):

a. find an equation for f^-1(x) b. graph f and f^-1 in the same rectangular coordinate system c. use interval notation to give the domain and the range of f and f^-1 f(x)=(x-1)^2,x_>1 f(x)=(x+2)^3

OpenStudy (anonymous):

\[f(x)=(x-1)^2,x \ge1\]

OpenStudy (turingtest):

\[y=(x-1)^2\]\[\sqrt y=x-1\]\[\sqrt y+1=x\]\[f^{-1}(x)=\sqrt x+1\]the domain is where f(x) is defined, so for this function\[[0,\infty)\]the range is the set of all possible values of the function, so since x must be at least zero the range is\[ [1,\infty)\]